# Federated Deep Reinforcement Learning for Privacy-Preserving Robotic-Assisted Surgery

### v1

@@ V Conclusion and Future Work
Experimental results validate the effectiveness of the FDRL framework in achieving an optimal balance between privacy and performance. The privacy-utility tradeoff analysis confirms that the framework successfully Minimises the PLR while preserving high surgical precision. Compared to centralized RL approaches, FDRL reduces the risk of data exposure, maintains model performance across diverse surgical tasks, and enhances policy generalization by leveraging institution-specific procedural knowledge. Additionally, policy divergence emerges as an implicit privacy-preserving measure, reducing the risk of reconstructing sensitive patient data from the shared global model.
Future work will clinically validate the robotic-assisted surgery framework with real patient data and partner hospitals, focusing on adaptability, generalization, and real-time performance under healthcare regulations. Key improvements target computational efficiency and latency using hardware-efficient strategies like quantised models, edge computing, and lightweight compression. In parallel, a critical direction will explore formal verification techniques to rigorously validate the correctness, safety, and privacy guarantees of the FL protocols and policy adaptation mechanisms. This includes employing model checking and formal methods to analyse decision-making sequences in high-assurance surgical settings, ensuring that the learned policies conform to medical safety constraints and privacy-preserving standards. Such verification methods will further strengthen the framework’s trustworthiness and clinical readiness, especially for regulatory approval in safety-critical applications. Additionally, security concerns related to privacy attacks, adversarial robustness, and reconstruction threats will be addressed. Scalability and personalization will improve through hospital-specific model fine-tuning, ensuring collaborative performance. The framework may also advance to real-time patient monitoring and remote diagnostics, emphasizing energy efficiency for underserved areas. This project advances privacy-preserving, adaptive, secure AI-driven robotic surgery, tackling key challenges in privacy, efficiency, and clinical integration.

### v2

@@ V Conclusion and Future Work
Experimental results validate the effectiveness of the FDRL framework in achieving an optimal balance between privacy and performance. The privacy-utility tradeoff analysis confirms that the framework successfully minimises the PLR while preserving high surgical precision. Compared to centralised RL approaches, FDRL reduces the risk of data exposure, maintains model performance across diverse surgical tasks, and enhances policy generalisation by leveraging institution-specific procedural knowledge. Additionally, policy divergence emerges as an implicit privacy-preserving measure, reducing the risk of reconstructing sensitive patient data from the shared global model.
Future work will clinically validate the robotic-assisted surgery framework with real patient data and partner hospitals, focusing on adaptability, generalisation, and real-time performance under healthcare regulations. Key improvements target computational efficiency and latency using hardware-efficient strategies like quantised models, edge computing, and lightweight compression. In parallel, a critical direction will explore formal verification techniques to rigorously validate the correctness, safety, and privacy guarantees of the FL protocols and policy adaptation mechanisms. This includes employing model checking and formal methods to analyse decision-making sequences in high-assurance surgical settings, ensuring that the learned policies conform to medical safety constraints and privacy-preserving standards. Such verification methods will further strengthen the framework’s trustworthiness and clinical readiness, especially for regulatory approval in safety-critical applications.
Additionally, security concerns related to privacy attacks, adversarial robustness, and reconstruction threats will be addressed. Scalability and personalisation will improve through hospital-specific model fine-tuning, ensuring collaborative performance. The framework may also advance to real-time patient monitoring and remote diagnostics, emphasizing energy efficiency for underserved areas. This project advances privacy-preserving, adaptive, secure AI-driven robotic surgery, tackling key challenges in privacy, efficiency, and clinical integration.